when the sum of 8 and 18 times a positive number is subtracted from the square of the number, zero results. Find the number?

w^2-(8w+18)=0
w^2-8w-18=0
It said to find the number so, would I factor out the binomial and set both to zero and find the number that way? I can't get it to factor out evenly. The closest I have got is (w-2)(w+9)=0
Can you please help me solve this

Answer by richwmiller(17219)

Not w^2-(8w+18)=0
but
w^2-(8+18w)=0
which still doesn't factor.

(w-9)^2-89 = 0
w = 9-sqrt(89) and w = 9+sqrt(89)
only the positive
w=18.434
no solution the way the problem is written.